# EXAONE Deep: Reasoning Enhanced Language Models

### v2

@@ 1 Introduction
Recently, there has been a growing trend in research to enhance reasoning performance by adjusting computing resources during the testing phase [16]. In response to this trend, LG AI Research is introducing a new model lineup called EXAONE Deep 2.4B, 7.8B, and 32B. These models are fine-tuned versions of the EXAONE 3.5 series [6], specifically optimized for reasoning tasks. We have trained these models using three prominent techniques widely employed for fine-tuning: Supervised Fine-Tuning (SFT), Direct Preference Optimization (DPO), and Online Reinforcement Learning (Online RL).
The performance evaluation of the models indicates that, for the 2.4B variant, it demonstrates superior performance compared to DeepSeek-R1-Distill-Qwen-1.5B [2]. For the 7.8B variant, it outperforms not only open-weight models of comparable scale such as DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B [2] but also a proprietary reasoning model OpenAI o1-mini [12]. In the case of the 32B model, its performance is competitive to that of the leading open-weight reasoning models such as QwQ-32B [13] and DeepSeek-R1 [2], and is superior to that of DeepSeek-R1-Distill-Qwen-32B and DeepSeek-R1-Distill-Llama-70B [2] as illustrated in Figure 1.
@@ 2.1 Data
SFT Dataset Example Query: Given the complex numbers $u$, $v$, and $w$ such that $uv=12+16i$, $vw=20-15i$, and $wu=-30+40i$, find real numbers $p$ and $q$ such that $u+v+w=p+qi$. Determine the value of $p^{2}+q^{2}$. Response: <thought> Okay, so I need to find real numbers $p$ and $q$ such that $u+v+w=p+qi$, and then compute $p^{2}+q^{2}$. The given information is the products of the complex numbers: $uv=12+16i$, $vw=20-15i$, and $wu=-30+40i$. Hmm, complex numbers can sometimes be tricky because of their real and imaginary parts, but maybe I can find expressions for $u$, $v$, $w$ individually? Let me think. If I have three equations involving the products of pairs of these complex numbers, maybe I can multiply all three equations together? Let’s see: $uv\times vw\times wu=(uvw)^{2}$. So $(uvw)^{2}$ = $(12+16i)(20-15i)(-30+40i)$. If I compute that product, I can then take the square root to find $uvw$. But taking the square root of a complex number might give multiple values, but since we’re dealing with complex numbers $u$, $v$, $w$, maybe we can choose a principal root or something. Wait, but maybe there’s a better way. Alternatively, if I can express each variable in terms of another. For example, from $uv=12+16i$, maybe $u=(12+16i)/v$. Similarly, from $vw=20-15i$, $w=(20-15i)/v$. Then from $wu=-30+40i$, substituting the expressions for $u$ and $w$ in terms of $v$, we get $[(20-15i)/v]\times[(12+16i)/v]=-30+40i$. That simplifies to $(20-15i)(12+16i)/(v^{2})=-30+40i$. Then solving for $v^{2}$, we get $v^{2}=(20-15i)(12+16i)/(-30+40i)$. Then once we find $v^{2}$, we can take square roots to find $v$, and subsequently find $u$ and $w$. …  (thought process continues) </thought> For each $v$, we compute $u$ and $w$: Case 1: $v=\frac{9-13i}{5}$ $u=-2+6i$ $w=\frac{15+5i}{2}$ Case 2: $v=\frac{-9+13i}{5}$ $u=2-6i$ $w=\frac{-15-5i}{2}$ Summing $u$, $v$, and $w$ in both cases: Case 1: $u+v+w=\frac{73}{10}+\frac{59}{10}i$ Case 2: $u+v+w=-\frac{73}{10}-\frac{59}{10}i$ In both cases, $p^{2}+q^{2}$ is: $\left(\frac{73}{10}\right)^{2}+\left(\frac{59}{10}\right)^{2}=\frac{73^{2}+59^{2}}{100}=\frac{5329+3481}{100}=\frac{8810}{100}=\frac{881}{10}$ Thus, the value of $p^{2}+q^{2}$ is: $\boxed{\dfrac{881}{10}}$
@@ 2.2 Training
The base models of EXAONE Deep are EXAONE 3.5 Instruct models [7, 9, 8], which are instruction-tuned models possessing instruction-following capabilities. To enhance the reasoning abilities of EXAONE Deep, we structured SFT and DPO data in a templated format as illustrated in Figure 3. Each training instance consists of a structured thought process followed by a final answer that synthesizes the reasoning steps into a coherent and precise response. Specifically, the EXAONE 3.5 Instruct models are trained to engage in reasoning within the <thought> and </thought> tags, performing step-by-step logical progression along with reflection, self-checking, and correction. The final answer, generated after reasoning, is designed to be self-contained, summarizing the key insights derived from the thought process in a clear and concise manner. This structured approach enables EXAONE Deep to engage in robust reasoning and deliver well-founded answers to a given query. We utilize SimPER [19] as the training algorithm for DPO and our designed GRPO [15] variant for Online RL.
@@ 3.1 Benchmarks
We evaluate models on MATH-500 [10], American Invitational Mathematics Examination (AIME) 2024 and 2025 [11], the mathematics sections of South Korea’s College Scholastic Ability Test (CSAT) 2025 [5], GPQA Diamond [14], LiveCodeBench (24.08-25.02) [4], MMLU [3], and MMLU-Pro [18].
@@ 3.3 Evaluation Setup
[…]
where $p_{i}$ denotes the correctness of the $i$-th response. We also report cons@$k$ [2] where the model generates $k$ answers and the final answer is the one generated the most frequently [17].
@@ 4 Limitations
The EXAONE Deep models introduced in this document are specifically fine-tuned to excel at reasoning tasks. Though their base models are instruction-fine-tuned and generally capable of following instructions, for addressing a wider range of real-world use cases, we strongly recommend utilizing the EXAONE 3.5 Instruct models [7, 9, 8], which are optimized for practical application scenarios.

### v3

@@ 1 Introduction
Recently, there has been a growing trend in research to enhance reasoning performance by adjusting computing resources during the testing phase [snell2024scalingllmtesttimecompute]. In response to this trend, LG AI Research is introducing a new model lineup called EXAONE Deep 2.4B, 7.8B, and 32B. These models are fine-tuned versions of the EXAONE 3.5 series [research2024exaone35serieslarge], specifically optimized for reasoning tasks. We have trained these models using three prominent techniques widely employed for fine-tuning: Supervised Fine-Tuning (SFT), Direct Preference Optimization (DPO), and Online Reinforcement Learning (Online RL).
The performance evaluation of the models indicates that, for the 2.4B variant, it demonstrates superior performance compared to DeepSeek-R1-Distill-Qwen-1.5B [deepseekai2025deepseekr1incentivizingreasoningcapability]. For the 7.8B variant, it outperforms not only open-weight models of comparable scale such as DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B [deepseekai2025deepseekr1incentivizingreasoningcapability] but also a proprietary reasoning model OpenAI o1-mini [openai2024o1systemcard]. In the case of the 32B model, its performance is competitive to that of the leading open-weight reasoning models such as QwQ-32B [qwen2025qwq32b] and DeepSeek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability], and is superior to that of DeepSeek-R1-Distill-Qwen-32B and DeepSeek-R1-Distill-Llama-70B [deepseekai2025deepseekr1incentivizingreasoningcapability] as illustrated in Figure 1.
@@ 2.1 Data
SFT Dataset Example Query: Given the complex numbers $u$, $v$, and $w$ such that $uv=12+16i$, vw=2015ivw=20-15i, and wu=30+40iwu=-30+40i, find real numbers $p$ and $q$ such that $u+v+w=p+qi$. Determine the value of $p^{2}+q^{2}$. Response: <thought> Okay, so I need to find real numbers $p$ and $q$ such that $u+v+w=p+qi$, and then compute $p^{2}+q^{2}$. The given information is the products of the complex numbers: $uv=12+16i$, vw=2015ivw=20-15i, and wu=30+40iwu=-30+40i. Hmm, complex numbers can sometimes be tricky because of their real and imaginary parts, but maybe I can find expressions for $u$, $v$, $w$ individually? Let me think. If I have three equations involving the products of pairs of these complex numbers, maybe I can multiply all three equations together? Let’s see: $uv\times vw\times wu=(uvw)^{2}$. So $(uvw)^{2}$ = (12+16i)(2015i)(30+40i)(12+16i)(20-15i)(-30+40i). If I compute that product, I can then take the square root to find uvwuvw. But taking the square root of a complex number might give multiple values, but since we’re dealing with complex numbers $u$, $v$, $w$, maybe we can choose a principal root or something. Wait, but maybe there’s a better way. Alternatively, if I can express each variable in terms of another. For example, from $uv=12+16i$, maybe $u=(12+16i)/v$. Similarly, from vw=2015ivw=20-15i, w=(2015i)/vw=(20-15i)/v. Then from wu=30+40iwu=-30+40i, substituting the expressions for $u$ and $w$ in terms of $v$, we get $(20-15i)/v]\times[(12+16i)/v]=-30+40i$. That simplifies to $(20-15i)(12+16i)/(v^{2})=-30+40i$. Then solving for $v^{2}$, we get $v^{2}=(20-15i)(12+16i)/(-30+40i)$. Then once we find $v^{2}$, we can take square roots to find $v$, and subsequently find $u$ and $w$. … (thought process continues) </thought> For each $v$, we compute $u$ and $w$: Case 1: $v=\frac{9-13i}{5}$ u=2+6iu=-2+6i $w=\frac{15+5i}{2}$ Case 2: $v=\frac{-9+13i}{5}$ u=26iu=2-6i $w=\frac{-15-5i}{2}$ Summing $u$, $v$, and $w$ in both cases: Case 1: $u+v+w=\frac{73}{10}+\frac{59}{10}i$ Case 2: $u+v+w=-\frac{73}{10}-\frac{59}{10}i$ In both cases, $p^{2}+q^{2}$ is: $\left(\frac{73}{10}\right)^{2}+\left(\frac{59}{10}\right)^{2}=\frac{73^{2}+59^{2}}{100}=\frac{5329+3481}{100}=\frac{8810}{100}=\frac{881}{10}$ Thus, the value of $p^{2}+q^{2}$ is: $\boxed{\dfrac{881}{10}}$
@@ 2.2 Training
The base models of EXAONE Deep are EXAONE 3.5 Instruct models [lgai2004exaone3524binstruct, lgai2004exaone3578binstruct, lgai2004exaone3532binstruct], which are instruction-tuned models possessing instruction-following capabilities. To enhance the reasoning abilities of EXAONE Deep, we structured SFT and DPO data in a templated format as illustrated in Figure 3. Each training instance consists of a structured thought process followed by a final answer that synthesizes the reasoning steps into a coherent and precise response. Specifically, the EXAONE 3.5 Instruct models are trained to engage in reasoning within the <thought> and </thought> tags, performing step-by-step logical progression along with reflection, self-checking, and correction. The final answer, generated after reasoning, is designed to be self-contained, summarizing the key insights derived from the thought process in a clear and concise manner. This structured approach enables EXAONE Deep to engage in robust reasoning and deliver well-founded answers to a given query. We utilize SimPER [xiao2025simperminimalistapproachpreference] as the training algorithm for DPO and our designed GRPO [shao2024deepseekmathpushinglimitsmathematical] variant for Online RL.
@@ 3.1 Benchmarks
We evaluate models on MATH-500 [lightman2023letsverifystepstep], American Invitational Mathematics Examination (AIME) 2024 and 2025 [maa2025aime], the mathematics sections of South Korea’s College Scholastic Ability Test (CSAT) 2025 [kice2025suneung], GPQA Diamond [rein2023gpqagraduatelevelgoogleproofqa], LiveCodeBench (24.08-25.02) [jain2024livecodebenchholisticcontaminationfree], MMLU [hendrycks2021measuringmassivemultitasklanguage], and MMLU-Pro [wang2024mmluprorobustchallengingmultitask].
@@ 3.3 Evaluation Setup
Following the setup described in the DeepSeek-R1 technical report [deepseekai2025deepseekr1incentivizingreasoningcapability], the maximum length of model generations is set to 32K tokens. Additionally, since the lengthy outputs of the reasoning models can vary greatly, as noted in the report, we adopt the pass@$k$ metric [chen2021evaluatinglargelanguagemodels] to ensure the reliability of model performances. Specifically, we generate $k$ responses for each test case using a sampling temperature of 0.6 and a top-$p$ value of 0.95000For QwQ-32B, we additionally applied the recommended $\text{top-}k=40$ during evaluation. For OpenAI o1-mini, we were unable to adjust sampling parameters to our desired settings; thus, we used temperature = 1 and $\text{top-}p=1$. , and calculate pass@1 as followed:
$\text{pass@1}=\frac{1}{k}\sum_{i=1}^{k}{p_{i}},$ (1)
where $p_{i}$ denotes the correctness of the $i$-th response. We also report cons@$k$ [deepseekai2025deepseekr1incentivizingreasoningcapability] where the model generates $k$ answers and the final answer is the one generated the most frequently [wang2023selfconsistencyimproveschainthought].
@@ 4 Limitations
The EXAONE Deep models introduced in this document are specifically fine-tuned to excel at reasoning tasks. Though their base models are instruction-fine-tuned and generally capable of following instructions, for addressing a wider range of real-world use cases, we strongly recommend utilizing the EXAONE 3.5 Instruct models [lgai2004exaone3524binstruct, lgai2004exaone3578binstruct, lgai2004exaone3532binstruct], which are optimized for practical application scenarios.